OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is parallel to the line x + 4y = 6 and passes through (-8, 5).

OpenStudy (anonymous):

first go ahead and put it into slope intercept form

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

how do i do that

OpenStudy (anonymous):

5=m(-8)+b

OpenStudy (anonymous):

agh sorry i dont think i remember enough. i know that the parallel line will have the same slope. then the y intercept will be your Y(im pretty sure.

OpenStudy (anonymous):

x+4y=6 is what you are given. put that into slope intercept form first.

OpenStudy (anonymous):

@Directrix could you help?

OpenStudy (anonymous):

do you know how to do that?

OpenStudy (anonymous):

not really i forget

OpenStudy (anonymous):

subtract the x from the one side and put it onto the other. what does that give you so far?

OpenStudy (anonymous):

4y=6x

OpenStudy (anonymous):

sorry I'm so tired and need to finish the last two problems on my worksheet

OpenStudy (anonymous):

no, you cant combine the 6 and x. so it would be 4y=-x+6. then you are trying to isolte the y. so divide by 4

OpenStudy (anonymous):

when someone done helping high, would you mind helping me?

OpenStudy (anonymous):

real quick

OpenStudy (anonymous):

y=-4x+1.5

OpenStudy (anonymous):

instead of 4x it would be -1/4x. do you know rise over run? or how to find slope?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

how long have you been working on stuff like this? its pretty important to know...hahah